OpenStudy (mewcifer):

I need help please An equation is shown below: 8(2x – 14) + 13 = 4x – 27 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points)

OpenStudy (anonymous):

http://www.geteasysolution.com/ Easy way to find your answer and it shows work! :D

OpenStudy (anonymous):

It's basically a calculator. :>

OpenStudy (mewcifer):

Thanks so much, do you think you can help me with another one? @StudyBuggTori

OpenStudy (anonymous):

Shoot! :> I'll try my best.

OpenStudy (mewcifer):

An equation is shown below: 8x + 2(x – 7) = 7x + 3x – 14 Part A: Solve the equation and write the number of solutions. Show all the steps. (6 points) Part B: Name one property you used to solve this equation. (4 points)

OpenStudy (mewcifer):

@StudyBuggTori

OpenStudy (anonymous):

You could use the same website for that one too. I know what I'm saying isn't much help but all your answers are there if you plug it in.

OpenStudy (mewcifer):

It's not working

OpenStudy (anonymous):

Simplifying 8x + 2(x + -7) = 7x + 3 + -14 Reorder the terms: 8x + 2(-7 + x) = 7x + 3 + -14 8x + (-7 * 2 + x * 2) = 7x + 3 + -14 8x + (-14 + 2x) = 7x + 3 + -14 Reorder the terms: -14 + 8x + 2x = 7x + 3 + -14 Combine like terms: 8x + 2x = 10x -14 + 10x = 7x + 3 + -14 Reorder the terms: -14 + 10x = 3 + -14 + 7x Combine like terms: 3 + -14 = -11 -14 + 10x = -11 + 7x Solving -14 + 10x = -11 + 7x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-7x' to each side of the equation. -14 + 10x + -7x = -11 + 7x + -7x Combine like terms: 10x + -7x = 3x -14 + 3x = -11 + 7x + -7x Combine like terms: 7x + -7x = 0 -14 + 3x = -11 + 0 -14 + 3x = -11 Add '14' to each side of the equation. -14 + 14 + 3x = -11 + 14 Combine like terms: -14 + 14 = 0 0 + 3x = -11 + 14 3x = -11 + 14 Combine like terms: -11 + 14 = 3 3x = 3 Divide each side by '3'. x = 1 Simplifying x = 1

OpenStudy (anonymous):

There you go. and here is the link for the first question. http://www.geteasysolution.com/8(2x-14)+13=4x-27